OpenStudy (anonymous):

Both triangles are similar. The area of the smaller triangle is about 97ft. Given Small triangle- 15ft Lager triangle -21ft

Directrix (directrix):

@barrazapris15 If two triangles are similar, then the square of the scale factor is equal to the ratio of the areas of the two triangles. (15/21)^2 = 95/x where x is the area of the larger triangle

OpenStudy (anonymous):

i got 133ft2

Directrix (directrix):

Just a sec. Let me see what I get.

Directrix (directrix):

4753/25 = 190.12 approx. I had a typo.

Directrix (directrix):

@barrazapris15